# Jasmine: Harnessing Diffusion Prior for Self-Supervised Depth Estimation

### v1

@@ 4.3 Ablation Study
Table 3: Ablation Studies. vK and Hy mean the virtual KITTI and Hypersim datasets. Dataset/$n$ denote we downsample the image to $\frac{1}{n}$ resloution(i.e. Hy/1.6 means downsample to 640$\times$192, where 640=1024/1.6) and resize them back.
@@
(c) Hybrid Image Reconstruction
(7) w/o HIR 0.132 0.673 4.271 0.852 0.967
@@
As shown in Table 3, we conduct ablation studies to validate our designs. Firstly, in sub-table (a), we gradually tested the effects of our basic components, such as SD prior, HIR, and SSG. The SD prior proves most critical - training from scratch (ID0 vs. ID1) causes catastrophic failure (AbsRel↑473%, RMSE↑206%). The other experiments also demonstrate that disabling HIR (ID4) or SSG (ID3) degrades performance by 47%/43% in AbsRel, proving the necessity of depth distribution alignment and SD detail preservation. In sub-table (b), we further melted down our proposed SSG. ID (6) reveals that the naive GRU-based depth residual refinement, which directly estimates the $D_{\delta}$ to update the depth, results in an inferior result. A comprehensive analysis of HIR was conducted in sub-table (c), where IDs (7-10, 0) correspond to Fig. 2 (b-f), respectively. From these experiments, we can draw similar conclusions to Sec. 3.2. The analysis of auxiliary images is presented in sub-table (d). The experiments in IDs (0, 13, 14) indicate that compared to real/synthesis datasets, the content of the dataset is more important, and diverse scenes offer greater benefits than street views similar to our primary dataset, KITTI. Furthermore, IDs (0, 15, 16) demonstrate that our surrogate task is robust to image sampling resolutions, as downsampling to $\frac{1}{1.6}$ or even $\frac{1}{4}$ has minimal impact on the results. Moreover, ID(13) further confirms that HIR remains effective even when trained on small-scale datasets (fewer than 1k samples). These insights provide potential opportunities for applying SD models to other dense estimation tasks and enhancing result sharpness.
In summary, these ablation results validate the effectiveness of our proposed adaptation protocol, indicating that each design plays a crucial role in optimizing the diffusion model for self-supervised depth estimation tasks.
@@ 5 Conclusion
We propose Jasmine, the first SD-based self-supervised framework for monocular depth estimation, effectively leveraging SD’s priors to enhance sharpness and generalization without high-precision supervision. To achieve this object, we introduce two novel modules: Hybrid Image Reconstruction (HIR) for mitigating reprojection artifacts and preserving the Stable Diffusion’s latent priors, alongside Scale-Shift GRU (SSG) to align scale-invariant depth predictions while suppressing noisy gradients. Extensive experiments demonstrate that Jasmine achieves SoTA performance on KITTI and superior zero-shot generalization across datasets. Our approach establishes a new paradigm for unsupervised depth estimation, paving the way for future advancements in self-supervised learning.
[…]
Supplementary Material
@@ Appendix A Proof of Scale-Invariant Depth
Note that for transformations within the same scene, the relative pose R remains constant[18]. Also, $a$ and $b$ here can be arbitrary numbers, so we have:
@@
$\zeta^{\prime}s_{c}D^{\prime}=K\left(\textbf{R}K^{-1}\zeta s_{c}D+s_{c}\textbf{T}\right)$ (17)
@@
Method Venue Notes Data Resolution AbsRel SqRel RMSE RMSElog $a_{1}$ $a_{2}$ $a_{3}$
ManyDepth[47] CVPR2021 ( -1,0) K(40K+0) 1024$\times$320 0.091 0.694 4.245 0.171 0.911 0.968 0.983
Dual Refine[1] CVPR2023 (-1,0) K(40K+0) 960$\times$288 0.087 0.674 4.130 0.167 0.915 0.969 0.984
Mono-ViFI[28] ECCV2024 (-1,0,1) K(40K+0) 1024$\times$320 0.089 0.556 3.981 0.164 0.914 0.971 0.986
EPCDepth[35] ICCV2021 Stereo K(45K+0) 1024$\times$320 0.091 0.646 4.207 0.176 0.901 0.966 0.983
PlaneDepth[45] CVPR2023 Stereo K(45K+0) 1280$\times$384 0.085 0.563 4.023 0.171 0.910 0.968 0.984
Jasmine - Mono KH(68K+0) 1024$\times$320 0.090 0.581 3.944 0.161 0.919 0.972 0.986
Appendix B Supervision Details
@@
SGD Loss for Dynamic Object.
@@
Building on this, we adopt the GDS loss (only the first stage) from[33] to handle dynamic objects. This approach is based on the smoothness loss and introduces a ground-contact-prior mask $M_{gr}$, defined as:
@@
where $M_{t}$ is the dynamic object segmentation obtained from a semantic segmentation network (1 represents dynamic and 0 for static), and $\gamma$ is the weighting parameter for $M_{gr}$, empirically set to 100. Considering the bottom pixels of dynamic regions like the car tire, they impose a high weighting on $|\partial_{y}\hat{d_{t}}|$ with $M_{gr}(i,j)=\gamma$, thereby enforcing its depth consists with its neighboring ground pixels below. So the final loss is:
@@
Edge Loss for Sharp Prediction.
We further introduce an edge loss to enhance the prediction’s details. Since the surrogate and primary tasks are decoupled after the UNet’s final layer, and the Scale-Shift Invariant (SSI) depth is shielded from photometric loss interference through SSG-based isolation, the SSI depth retains significantly richer structural details. To transfer these details to the final result, we introduce a simple edge loss. Specifically, we design a GradNet,
@@
Sky Loss for Anti-artifact.
@@
Eliminating the Bottleneck of Teacher Loss
@@
Furthermore, to avoid the performance ceiling imposed by the teacher model, we draw inspiration from D4RD[44], and employ an adaptive filtering mechanism:
@@
The segmentation model mentioned above is Mask2Former[6] throughout. To avoid the misconception that the details stem from the segmentation network’s output, we further expand Fig. 2 (in paper) with segment-based methods. In fact, to preserve these details, a naive approach is to employ semantic segmentation constraints[33]. However, as shown in Fig. 7 (a-2), using the semantic triplet loss from [33] not only disrupts the depth distribution but also introduces spurious edges, rendering it incompatible with SD’s latent priors. Furthermore, the results in paper Table 1 are compared with a segmentation-based model Jaeon et al*, and our method achieves superior performance.
Appendix C Evaluation Metrics
Similar to [15], we employ the following evaluation metrics in our experiments,
AbsRel: $\ \frac{1}{|{M}_{vl}|}\sum_{d\in{M}_{vl}}\left|d-d_{gt}\right|/d_{gt}$;
RMSE: $\sqrt{\frac{1}{|{M}_{vl}|}\sum_{d\in{M}_{vl}}\left\|d-d_{gt}\right\|^{2}}$;
${a}_{t}$: percentage of $d$ such that $\max(\frac{d}{d_{gt}},\frac{d_{gt}}{d})<1.25^{t}$ ; where $d_{gt}$ and $d$ denote the GT and estimated pixel depth, ${M}_{vl}$ is the valid mask set to $801e-3<d_{gt}<$ 80.
Appendix D Addition Experiments Results
Regarding SQLdepth
SQLdepth’s authors clarified that they used a new KITTI training split, which included more data than the commonly used split, totaling 70k samples. Later, it was discovered that this training set had a risk of overlapping with the test set, as detailed in the repository’s issues. Consequently, we chose not to compare our work with SQLdepth or its subsequent studies.
Comparison with Other Self-Supervised Settings
@@ Appendix D Addition Experiments Results
As shown in Table 4, although the multi-frame and stereo-supervision models leverage more information during training and inference, our method still outperforms the state-of-the-art approaches in the corresponding domains across most metrics. This further demonstrates the superiority of our approach.
Hybird Ratio Ablation Details

### v2

@@ 4.4 Ablation Study
As shown in Table 3, we conduct ablation studies to validate our designs. Firstly, in sub-table (a), we gradually tested the effects of our basic components, such as SD prior, MIR, and SSG. The SD prior proves most critical - training from scratch (ID0 vs. ID1) causes catastrophic failure (AbsRel↑473%, RMSE↑206%). The other experiments also demonstrate that disabling MIR (ID4) or SSG (ID3) degrades performance by 47%/43% in AbsRel, proving the necessity of depth distribution alignment and SD detail preservation.
[…]
@@
(c) Mix-batch Image Reconstruction
(7) w/o MIR 0.132 0.673 4.271 0.852 0.967
@@
In sub-table (b), we further melted down our proposed SSG. ID (6) reveals that the naive GRU can initially solve the distribution misalignment by estimating the $D_{\delta}$ (Eq. 4). However, the scale difference between SSI and SI depth makes it difficult to restore through linear addition. Therefore, after introducing SST, the overall model performance is further enhanced by 10%, ultimately achieving SoTA performance. A comprehensive analysis of MIR was conducted in sub-table (c) and we can draw similar conclusions to Sec. 3.2. Jasmine significantly outperforms the alternatives, such as KITTI/synthetic-only reconstruction (IDs (8,9) and Fig. 3 (c,d)) and latent-space supervision (ID (10) and Fig. 3 (e)), which strongly proves the effectiveness of our proposed MIR. The analysis of auxiliary images is presented in sub-table (d). The experiments in IDs (0, 13, 14) indicate that, compared to real/synthesis datasets, the content of the dataset is more important, and diverse scenes offer greater benefits than street views (virtual KITTI images) similar to our primary dataset, KITTI. Furthermore, IDs (0, 15, 16) demonstrate that our surrogate task is robust to image sampling resolutions, as downsampling to $\frac{1}{1.6}$ or even $\frac{1}{4}$ has minimal impact on the results. Moreover, ID(13) further confirms that MIR remains effective even when trained on small-scale datasets (fewer than 1k samples). These insights provide potential opportunities for applying SD models to other dense estimation tasks and enhancing result sharpness. In summary, these ablation results validate the effectiveness of our proposed adaptation protocol, indicating that each design plays a crucial role in optimizing the diffusion model for self-supervised depth estimation tasks.
4.5 Inference Latency
[…]
Method Marigold Lotus E2E-Mono Monodepth2 MonoViT MonoViFi Jasmine
MACs 133T 2.65T 2.65T 21.43G 25.63G 28.79G 2.83T
Runtime 9.88s 157ms 152ms 33ms 29ms 25ms 172ms
@@ 5 Conclusion
We propose Jasmine, the first SD-based self-supervised framework for monocular depth estimation, effectively leveraging SD’s priors to enhance sharpness and generalization without high-precision supervision. To achieve this objective, we introduce two novel modules: Mix-batch Image Reconstruction (MIR) for mitigating reprojection artifacts and preserving Stable Diffusion’s latent priors, alongside Scale-Shift GRU (SSG) to align scale-invariant depth predictions while suppressing noisy gradients. Extensive experiments demonstrate that Jasmine achieves SoTA performance on KITTI and superior zero-shot generalization across datasets. Our approach establishes a new paradigm for unsupervised depth estimation, paving the way for future advancements in self-supervised learning.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant (U2441242,62172032) and Graduate Research Innovation Project under Grant (KKYJS25001536).
Appendix
In this appendix, we provide more implementation details, experiments, analysis, and discussions for a comprehensive evaluation and understanding of Jasmine. Detailed contents are listed as follows:
@@ Appendix A Proof of Scale-Invariant Depth
Note that for transformations within the same scene, the relative pose R remains constantssi2si . Also, $a$ and $b$ here can be arbitrary numbers, so we have:
@@
$\zeta^{\prime}s_{c}D^{\prime}=K\left(\textbf{R}K^{-1}\zeta s_{c}D+s_{c}\textbf{T}\right),$ (15)
@@
Method Venue Notes Data AbsRel SqRel RMSE RMSElog $a_{1}$ $a_{2}$ $a_{3}$
ManyDepthmanydepth CVPR2021 ( -1,0) K(40K+0) 0.091 0.694 4.245 0.171 0.911 0.968 0.983
Dual Refinedualrefine CVPR2023 (-1,0) K(40K+0) 0.087 0.674 4.130 0.167 0.915 0.969 0.984
Mono-ViFIvifi ECCV2024 (-1,0,1) K(40K+0) 0.089 0.556 3.981 0.164 0.914 0.971 0.986
EPCDepthepcdepth ICCV2021 Stereo K(45K+0) 0.091 0.646 4.207 0.176 0.901 0.966 0.983
PlaneDepthplanedepth CVPR2023 Stereo K(45K+0) 0.085 0.563 4.023 0.171 0.910 0.968 0.984
Jasmine - Mono KH(68K+0) 0.090 0.581 3.944 0.161 0.919 0.972 0.986
Appendix B Detailed Training Workflow and Pseudocode
B.1 Step-by-Step Workflow
In this section, we temporarily omit the specifics of MIR and SSG to clarify the integration of self-supervision with the Stable Diffusion framework in Jasmine:
  • Input: A sequence of temporally adjacent images (like video frames), e.g., a source image $I_{t^{\prime}}$ and a target image $I_{t}$.
    • Here $I_{t}$ is exactly the training image, $I_{t^{\prime}}$ can be the previous frame $I_{t-1}$ or the next frame $I_{t+1}$ of $I_{t}$.
  • Step 1: Depth Prediction (via SD U-Net):
    • The SD U-Net, $f_{\theta}^{z}$, takes the latent of the target image, $z^{I}=\text{VAE.encode}(I_{t})$, and a pure noise vector, $\mathbf{n}$, as input.
    • It performs a single-step denoising process to predict the latent representation of a depth map, $z_{0}^{y}$.
        This is the key point we mentioned in Sec 3.1: single-step denoising make this step become a fast, feed-forward process rather than computationally prohibitive with slow, iterative denoising.
    • The VAE decoder then converts $z_{0}^{y}$ into the final depth map, $D=\text{VAE.decode}(z_{0}^{y})$.
    • Crucially, no GT depth is ever used in this step.
  • Step 2: Pose Prediction:
    • A separate PoseNet takes the image pair ($I_{t}$, $I_{t^{\prime}}$) as input and predicts the relative camera pose (rotation and translation), $T_{t\to t^{\prime}}$.
  • Step 3: Self-Supervised Signal Generation (Image Reprojection):
    • Using the predicted depth map $D$ and the predicted pose $T_{t\to t^{\prime}}$, we perform a warping operation to reproject the pixels from the source image $I_{t^{\prime}}$ onto the target image’s coordinate system. This creates a synthesized target image, $\hat{I}_{t}$.
  • Step 4: Loss Calculation and Optimization:
    • A photometric reprojection loss, $L_{ph}$, is calculated by comparing the synthesized image $\hat{I}_{t}$ with the original target image $I_{t}$.
    • This loss, $L_{ph}$, is the core self-supervised signal. If the predicted depth $D$ is incorrect, the reprojected image $\hat{I}_{t}$ will not match the original $I_{t}$, resulting in a high loss.
  • Step 5: End-to-End Backpropagation:
    • The gradient from $L_{ph}$ is backpropagated through the entire computational graph. This means the gradient flows back to update the weights of both the PoseNet and, most importantly, the SD U-Net ($f_{\theta}^{z}$).
  • Output: The predicted depth map $D$ obtained from Step 1.
B.2 Training Pseudocode
Algorithm 1 outlines the complete training procedure for Jasmine, incorporating all components including MIR, SSG, and the full loss computation (Corresponds to the pipeline in Fig. 2).
Algorithm 1 Jasmine Training Algorithm
1:Initialize: SD U-Net, PoseNet, SSG, VAE, optimizer
2:VAE.eval() $\triangleright$ VAE weights are frozen
3:for each batch of inputs ‘I’ in dataloader do
4:  $\triangleright$ Mix-batch Image Reconstruction (MIR)
5:  $s_{depth}\leftarrow[1,0]$, $s_{recon}\leftarrow[0,1]$ $\triangleright$ Task switchers
6:  $z^{I}_{t},z^{I}_{m},z^{n}\leftarrow\text{VAE.encode}(\text{I['I\_t']},\text{I['I\_m']},\text{noise})$
7:  $D_{ssi}\leftarrow\text{VAE.decode}(\text{SD\_UNet}([z^{I}_{t},z^{n}],t=999,s=s_{depth}))$
8:  $I_{rec}\leftarrow\text{VAE.decode}(\text{SD\_UNet}([z^{I}_{m},z^{n}],t=999,s=s_{recon}))$
9:  $\triangleright$ Scale-Shift GRU (SSG)
10:  $D_{list}\leftarrow\text{SSG}(D_{ssi},z^{I}_{t})$ $\triangleright$ $D_{list}$ contains $D_{ssi},D_{1},D_{si}]$
11:  $pose\leftarrow\text{PoseNet}(\text{I['I\_t']},\text{I['I\_t'']})$ $\triangleright$ Pose Estimation for Self-Supervision
12:  $loss\leftarrow\text{compute\_loss}(\text{I},D_{list},pose,I_{rec})$ $\triangleright$ Loss Computation and Optimization
13:  optimizer.zero_grad()
14:  lossloss.backward()
15:  optimizer.step()
16:end for
17:function compute_loss(inputs, $D_{list}$, pose, $I_{rec}$)
18:  $D_{si},D_{ssi}\leftarrow D_{list}[-1],D_{list}[0]$
19:  $reproj\_img\leftarrow\text{reproject}(\text{inputs['I\_t'']},D_{si},pose)$
20:  $L_{ph}\leftarrow\text{photometric\_loss}(reproj\_img,\text{inputs['I\_t']})$ $\triangleright$ Core self-supervised signal
21:  $L_{s}\leftarrow\text{photometric\_loss}(I_{rec},\text{inputs['I\_m']})$ $\triangleright$ Surrogate task signal
22:  $L_{tc}\leftarrow\text{teacher\_loss}(D_{list},\text{inputs['D\_tc']})$ $\triangleright$ As per Eq. 11
23:  $L_{a}\leftarrow\text{compute\_auxiliary\_loss}(\text{inputs},D_{si},D_{ssi})$
24:  $loss\leftarrow L_{ph}+L_{s}+L_{tc}+L_{a}\cdot 0.008$
25:  return lossloss
26:end function
27:function compute_auxiliary_loss(inputs, $D_{si},D_{ssi}$)
28:  $L_{GDS}\leftarrow\text{gds\_loss}(\text{inputs['I\_t']},\text{inputs['seg']},D_{si})$
29:  $L_{SKY}\leftarrow\text{sky\_loss}(D_{ssi},\text{inputs['sky\_mask']})$
30:  $L_{e}\leftarrow\text{edge\_loss}(D_{si},D_{ssi})$
31:  return $L_{GDS}+L_{SKY}+L_{e}$
32:end function
Appendix C Supervision Details
@@
C.1 SGD Loss for Dynamic Object
@@
Building on this, we adopt the GDS loss (only the first stage) fromgds to handle dynamic objects. This approach is based on the smoothness loss and introduces a ground-contact-prior mask $M_{gr}$, defined as:
@@
[…]
@@
C.2 Edge Loss for Sharp Prediction
We further introduce an edge loss to enhance the prediction’s details. Since the surrogate and primary tasks are decoupled after the U-Net’s final layer, and the Scale-Shift Invariant (SSI) depth is shielded from photometric loss interference through SSG-based isolation, the SSI depth retains significantly richer structural details. To transfer these details to the final result, we introduce a simple edge loss. Specifically, we design a GradNet,
@@ C.2 Edge Loss for Sharp Prediction
C.3 Sky Loss for Anti-artifact
@@
C.4 Eliminating the Bottleneck of Teacher Loss
@@
Furthermore, to avoid the performance ceiling imposed by the teacher model, we draw inspiration from D4RDd4rd , and employ an adaptive filtering mechanism:
@@
The segmentation model mentioned above is Mask2Formermask2former throughout. To avoid the misconception that the details stem from the segmentation network’s output, we further expand Fig. 3 (in paper) with segment-based methods. In fact, to preserve these details, a naive approach is to employ semantic segmentation constraintsgds . However, as shown in Fig. 8 (a-2), using the semantic triplet loss from gds not only disrupts the depth distribution but also introduces spurious edges, rendering it incompatible with SD’s latent priors. Furthermore, the results in paper Table 1 are compared with a segmentation-based model Jaeon et al*, and our method achieves superior performance.
Appendix D Preliminaries of GRU
The Gated Recurrent Unit (GRU) is a type of recurrent neural network (RNN) that uses gating mechanisms to control the flow of information between the previous hidden state and the current input, making it effective for sequence modeling tasks. Recently, GRUs have been gradually used in depth estimationiebins ; nddepth . A GRU cell updates its hidden state $\mathbf{h}^{k+1}$ based on the previous hidden state $\mathbf{h}^{k}$ and the current input $\mathbf{x}^{k}$. This update process is primarily managed by two gates: the reset gate ($\mathbf{r}$) and the update gate ($\mathbf{z}$). The reset gate determines how much of the past information (from the previous hidden state) to effectively "forget" or "reset" before computing a new candidate hidden state. The update gate then decides how much of this new candidate hidden state should be incorporated into the final hidden state, versus how much of the previous hidden state to carry over. The mathematical formulation for these operations is as follows:
$\begin{array}[]{l}\mathbf{z}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{z}),\quad\mathbf{r}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{r}),\\ \widehat{\mathbf{h}}^{k+1}=\tanh({Conv}([\mathbf{r}^{k+1}\odot\mathbf{h}^{k},\mathbf{x}^{k}],W_{h})),\\ \mathbf{h}^{k+1}=(1-\mathbf{z}^{k+1})\odot\mathbf{h}^{k}+\mathbf{z}^{k+1}\odot\widehat{\mathbf{h}}^{k+1},\end{array}$ (23)
where $k$ denotes iteration steps, $\sigma$ is the sigmoid activation function (outputting values between 0 and 1, ideal for gating), and $\odot$ indicates element-wise multiplication. The terms $W_{z},W_{r},W_{h}$ represent the learnable parameters (typically weight matrices and biases) for the update gate, reset gate, and candidate hidden state computation, respectively. The Conv{Conv} notation suggests that convolutional layers are used for these transformations, as is common when applying GRUs to feature maps in computer vision tasks. In our paper, the refined hidden state $\mathbf{h}^{k+1}$ predicts depth adjustment $D_{\delta}$ to update $D_{k}$, yielding $D_{k+1}$ for subsequent iterations, as utilized in our Scale-Shift GRU (SSG) module (see Sec. 3.3).
Appendix E Addition Experiments Results
E.1 Evaluation on KITTI Improved Benchmark
The standard self-supervised evaluation on the KITTI dataset is typically conducted using the raw LiDAR GT. However, due to the noisy nature of LiDAR data and known preprocessing issues [^1] Discussed in MonoDepth2 repository: https://github.com/nianticlabs/monodepth2/issues/274, we also provide the eigen improved benchmark result. Following the practice adopted in Mono-ViFi vifi , as shown in Table 5, this evaluation further confirms Jasmine’s SoTA performance, where it consistently outperforms other leading methods.
Table 5: Quantitative results on the KITTI dataset using the improved GT impr . The performance of SD-based methods is not fully reported in their original papers.
[…]
Marigold marigold 0.099 - - - 0.916 0.987 -
E2E-FT e2eft 0.096 - - - 0.921 0.980 -
Lotus lotus 0.081 - - - 0.931 0.987 -
Jasmine* 0.064 0.294 2.982 0.097 0.957 0.994 0.998
MonoViT monovit 0.068 0.314 3.125 0.105 0.948 0.992 0.998
MonoViFi vifi 0.071 0.338 3.539 0.113 0.937 0.990 0.998
Jasmine 0.061 0.255 2.765 0.092 0.963 0.995 0.999
E.2 Comparison with Other Self-Supervised Settings
As mentioned in Sec. 2, our single-frame monocular approach is more challenged but practical compared to other self-supervised configurations. Compared to stereo-based methods, monocular methods face additional challenges with dynamic objects and pose estimation inaccuracy, but monocular methods can eliminate the need for synchronized binocular cameras and precise calibration. Similarly, single-frame models neglect temporal information during inference, while multi-frame methods leverage consecutive frames to construct cost volumes and even support iterative refinement at test time, yielding improved accuracy. However, the practical deployment of multi-frame methods remains constrained by the availability of multiple frames.
@@ E.2 Comparison with Other Self-Supervised Settings
Despite these inherent disadvantages, as shown in Table 4, our method still outperforms the state-of-the-art approaches in both multi-frame and stereo-supervision domains across most metrics. This remarkable achievement, especially considering our more challenging problem setting, further demonstrates the substantial strength and generalizability of our approach.
E.3 Qualitative Comparisons